If Yao Ming, the 7-foot 6-inch center of the Houston Rockets casts a shadow of 18 feet 9 inches and Chauncy Billups guard with the Detroit Pistons, standing beside him casts a shadow of 15 feet, 7 inches, how tall is Billups?
Answer by DrBeeee(684)

Let's avoid decimals by converting all measurments to inches.
Yoa is 7'-6" or (12*7 + 6)" or 90"
Yao's shadow is 18'-9" or (12*18+9)" or 225"
Chauncy's shadow is 15'-7" or (12*15+7)" or 187"
Let C = Chauncy's height.
Use the height/shadow proportion
(1) C/187 = 90/225 or
(2) C = (187*90)/225 or
(3) C = 74.8 or
(4) C = 6'-2.8"
Answer: Chauncy Billups is 6' and 2.8" tall